# ReFoRCE: A Text-to-SQL Agent with Self-Refinement, Format Restriction, and Column Exploration

### v2

@@ Abstract
Text-to-SQL systems have unlocked easier access to critical data insights by enabling natural language queries over structured databases. However, deploying such systems in enterprise environments remains challenging due to factors such as large, complex schemas ( columns), diverse SQL dialects (e.g., BigQuery, Snowflake) and sophisticated query requirements (e.g., transformation, analytics). Current state-of-the-art performance on the Spider 2.0 dataset — a benchmark built to mimic such complex environments — remains limited at 20%. Key limitations include inadequate instruction-following, poor long-context comprehension, weak self-refinement, and insufficient dialect-specific knowledge. To address these gaps, we propose ReFoRCE (Self-Refinement Agent with Format Restriction and Column Exploration) which introduces (1) table compression to mitigate long-context limitations (2) format restriction to ensure accurate answer format, and (3) iterative column exploration for enhanced schema understanding. Additionally, it employs self-refinement pipeline consisting of (1) parallelized workflows with voting mechanisms and (2) a Common Table Expression (CTE) based refinement approach to handle unresolved cases. ReFoRCE achieves state-of-the-art results scoring 31.26 on the Spider 2.0-Snow and scoring 30.35 on the Spider 2.0-Lite tasks.
@@ 1 Introduction
Previous Text-to-SQL research primarily focused on model training and fine-tuning (Zhong et al., 2017; Wang et al., 2019; Scholak et al., 2021) on simpler datasets like Spider 1.0 (Yu et al., 2018). The rise of large language models (LLMs) has shifted this paradigm from ”pre-train and fine-tune” to prompting, thanks to the strong code generation capabilities of LLMs (Anthropic, 2023; Roziere et al., 2023; Achiam et al., 2023). Consequently, numerous Text-to-SQL methods now rely on prompting (Zhang et al., 2023; Gao et al., 2023a; Pourreza & Rafiei, 2024; Talaei et al., 2024) with powerful LLM API services.
@@ 3.1 Preliminaries
Lei et al. (2024) introduced Spider-Agent, a framework built on the ReAct Yao et al. (2023) paradigm with function-calling capabilities such as EXEC_SQL for executing SQL queries and TERMINAL for performing command-line operations to navigate DBT (Data Build Tool) projects and read schema-related files. The agent operates by receiving observations, which represent the current state of the environment or the outcome of a function call initiated by the agent. Based on these observations, the agent generates a ”Thought” and selects an appropriate ”Action” from a predefined list of function calls. The task is considered complete when the agent invokes the TERMINATE function.
@@ 3.2.1 Table Information Compression
Following the approach of Spider 2.0-Snow (Lei et al., 2024), we create a dictionary for each example, incorporating external knowledge and table structures using Database Definition Language (DDL) files. In some specific examples, DDL files exceed 300 KB, surpassing the context limitations of models like ChatGPT. To address this, we apply a pattern-based matching that merge tables with similar prefixes or suffixes. For these tables, we retain only one representative DDL file as input, while for others, we provide only the table names to the model.
@@ 3.2.2 Expected Answer Format Restriction
To address this, we propose Expected Answer Format Restriction, which involves generating the expected format at the outset and consistently reinforcing this format during self-refinement. The response must strictly adhere to the specified format in CSV style, ensuring alignment with executed CSV files. Each column should be explicitly defined, including all necessary attributes, and each record should occupy a separate row. The format should account for specific cases, such as superlatives, percentages, or coordinates, ensuring the output is concise, clear, and unambiguous. For ambiguous terms, potential values or additional columns can be added to maintain clarity and precision. Figure 1 illustrates a case of format restriction. For example, when a question asks for the highest number, the answer should be presented in a single row. Annotations such as ”answer in one row” can be appended to the format table to ensure clarity. Additionally, in the Spider 2.0 evaluation setting (Lei et al., 2024), it is acceptable to include extra columns. For instance, even if the task only requires barcodes, our format also includes a ”copy number” column, which is highly relevant to the task. This approach ensures consistency and accuracy, even when managing long contexts and complex task descriptions.
@@ 3.2.3 Exploration of Potentially Useful Columns
To address these challenges and ensure a comprehensive understanding of the database structure, we design a systematic approach to explore potentially useful columns. The process begins with identifying relevant tables and columns, guided by prompts designed to extract meaningful information. Dynamically generated SQL queries progress from simple, non-nested formats to more complex ones, enabling a gradual understanding of the database and arriving at the correct answer. These queries follow the structure SELECT DISTINCT "COLUMN_NAME" FROM DATABASE.SCHEMA.TABLE WHERE ... (specific to the Snowflake dialect) and adhere to constraints such as avoiding Common Table Expressions (CTEs) and schema-level checks, while limiting the output to 100 rows or 5 KB per query. Columns in JSON or nested formats are explicitly handled using techniques like LATERAL FLATTEN to extract nested values. Additionally, string-matching queries utilize fuzzy patterns (e.g., %target_str%) to enhance flexibility.
@@ Appendix A Algorithm for SQLs List Execution with Self-Correction during Column Exploration
The algorithm describes a process for executing a list of SQL queries with a self-correction mechanism during column exploration. It iteratively processes each SQL query from the input list, executes it, and evaluates the results. If the results are valid and complete, the query and its output are stored. If the results indicate issues (e.g., missing or invalid data), the algorithm attempts to correct the query using a self-correction method that generates and refines SQL queries based on error feedback, up to a maximum of three iterations. During this process, additional corrections for similar queries are dynamically generated by interacting with a large language model (LLM) chat session. The model suggests updated SQL queries, which can replace the remaining queries in the list if they are deemed sufficient. This iterative approach ensures robustness in handling invalid results and improves query accuracy by dynamically adapting to errors, while maintaining a focus on generating meaningful results for the task.
@@ Appendix B Algorithm for Self-Refinement Workflow for SQL Query Execution
Check if nested values.
@@
The algorithm describes a self-refinement workflow for iteratively generating, executing, and refining SQL queries. It begins by constructing an initial query environment () that incorporates input prompts, including table information, task descriptions, column exploration strategies, format restrictions, and API-specific query guidelines. During each iteration, a new SQL query is generated using the chat session model and executed via the specified API. If the query returns valid results, the results are processed by rounding numeric columns to two decimal places and handling nested values. The results are checked for uniqueness and format compliance, and if they satisfy these criteria, they are stored in a results table (results_tables). Self-consistency is achieved by verifying if the same result appears in results_tables twice, at which point the process terminates. If errors or invalid data are encountered, the algorithm refines the query by addressing issues such as missing or empty columns, improper handling of nested data, or syntax errors, and retries the process with the updated query. The workflow terminates either when valid results meet self-consistency, when the maximum number of iterations (max_iter) is reached, or when three consecutive errors occur, ensuring incomplete or invalid outputs are discarded. This structured approach improves robustness and accuracy in SQL query execution.

### v3

@@ Abstract
Text-to-SQL systems have unlocked easier access to critical data insights by enabling natural language queries over structured databases. However, deploying such systems in enterprise environments remains challenging due to factors such as large, complex schemas ( columns), diverse SQL dialects (e.g., BigQuery, Snowflake) and sophisticated query requirements (e.g., transformation, analytics). Current state-of-the-art performance on the Spider 2.0 dataset — a benchmark built to mimic such complex environments — remains limited at 20%. Key limitations include inadequate instruction-following, poor long-context comprehension, weak self-refinement, and insufficient dialect-specific knowledge. To address these gaps, we propose ReFoRCE (Self-Refinement Agent with Format Restriction and Column Exploration) which introduces (1) table compression to mitigate long-context limitations (2) format restriction to ensure accurate answer format, and (3) iterative column exploration for enhanced schema understanding. Additionally, it employs self-refinement pipeline consisting of (1) parallelized workflows with voting mechanisms and (2) a Common Table Expression (CTE) based refinement approach to handle unresolved cases. ReFoRCE achieves state-of-the-art results scoring 31.26 on the Spider 2.0-Snow and scoring 30.35 on the Spider 2.0-Lite tasks. Our code is available at https://github.com/hao-ai-lab/ReFoRCE.
@@ 1 Introduction
Previous Text-to-SQL research primarily focused on model training and fine-tuning (Zhong et al., 2017; Wang et al., 2019; Scholak et al., 2021) on simpler datasets like Spider 1.0 (Yu et al., 2018). The rise of large language models (LLMs) has shifted this paradigm from “pre-train and fine-tune” to prompting, thanks to the strong code generation capabilities of LLMs (Anthropic, 2023; Roziere et al., 2023; Achiam et al., 2023). Consequently, numerous Text-to-SQL methods now rely on prompting (Zhang et al., 2023; Gao et al., 2023a; Pourreza & Rafiei, 2024; Talaei et al., 2024) with powerful LLM API services.
@@ 3.1 Preliminaries
Lei et al. (2024) introduced Spider-Agent, a framework built on the ReAct Yao et al. (2023) paradigm with function-calling capabilities such as EXEC_SQL for executing SQL queries and TERMINAL for performing command-line operations to navigate DBT (Data Build Tool) projects and read schema-related files. The agent operates by receiving observations, which represent the current state of the environment or the outcome of a function call initiated by the agent. Based on these observations, the agent generates a “Thought” and selects an appropriate “Action” from a predefined list of function calls. The task is considered complete when the agent invokes the TERMINATE function.
@@ 3.2.1 Table Information Compression
Following the approach of Spider 2.0 (Lei et al., 2024), we create a dictionary for each example, incorporating external knowledge and table structures using Database Definition Language (DDL) files. In some specific examples, DDL files exceed setting size, surpassing the context limitations of models like ChatGPT. To address this, we apply a pattern-based matching that merge tables with similar prefixes or suffixes. For these tables, we retain only one representative DDL file as input, while for others, we provide only the table names to the model. Additionally, we use a straightforward API-based schema linking approach for examples with relatively long prompts to ensure the context length remains within limits while minimizing information loss during the linking process.
@@ 3.2.2 Expected Answer Format Restriction
To address this, we propose Expected Answer Format Restriction, which involves generating the expected format at the outset and consistently reinforcing this format during self-refinement. The response must strictly adhere to the specified format in CSV style, ensuring alignment with executed CSV files. Each column should be explicitly defined, including all necessary attributes, and each record should occupy a separate row. The format should account for specific cases, such as superlatives, percentages, or coordinates, ensuring the output is concise, clear, and unambiguous. For ambiguous terms, potential values or additional columns can be added to maintain clarity and precision. Figure 1 illustrates a case of format restriction. For example, when a question asks for the highest number, the answer should be presented in a single row. Annotations such as “answer in one row” can be appended to the format table to ensure clarity. Additionally, in the Spider 2.0 evaluation setting (Lei et al., 2024), it is acceptable to include extra columns. For instance, even if the task only requires barcodes, our format also includes a “copy number” column, which is highly relevant to the task. This approach ensures consistency and accuracy, even when managing long contexts and complex task descriptions.
@@ 3.2.3 Exploration of Potentially Useful Columns
To address these challenges and ensure a comprehensive understanding of the database structure, we design a systematic approach to explore potentially useful columns. The process begins with identifying relevant tables and columns, guided by prompts designed to extract meaningful information. Dynamically generated SQL queries progress from simple, non-nested formats to more complex ones, enabling a gradual understanding of the database and arriving at the correct answer. These queries follow the structure SELECT DISTINCT "COLUMN_NAME" FROM DATABASE.SCHEMA.TABLE WHERE ... (specific to the Snowflake dialect) and adhere to constraints such as avoiding Common Table Expressions (CTEs) and schema-level checks, while limiting the output to 100 rows or 500 B per query. Columns in JSON or nested formats are explicitly handled using techniques like LATERAL FLATTEN to extract nested values. Additionally, string-matching queries utilize fuzzy patterns (e.g., %target_str%) to enhance flexibility.
@@ Appendix A Algorithm for SQLs List Execution with Self-Correction during Column Exploration
Algorithm 1 begins by initializing an empty dictionary to store query results and a list to record errors. It then processes a list of SQL queries sequentially by executing each query and immediately checking the validity of its results. If a query returns an invalid result, the algorithm stores a corrected version of the query along with its result. Otherwise, it enters a self-correction loop where, for a limited number of iterations, it attempts to refine the query by generating a corrected version using a self-correction function and re-executes it. Once a corrected query produces valid results, these are added to the result dictionary. Additionally, if the successful result meets certain conditions (being a non-empty valid string), the algorithm logs this event and, if the error frequency suggests a systemic issue, it replaces a significant portion of the remaining queries with a new set of SQL queries generated from a correction prompt. This iterative process continues until all queries are processed or an excessive number of errors prompts an early termination of the execution loop.
@@ Appendix B Algorithm for Self-Refinement Workflow for SQL Query Execution
Check if nested values, null columns.
@@
Algorithm 2 describes a self-refinement workflow for iteratively generating, executing, and refining SQL queries. It begins by constructing an initial query environment () that incorporates input prompts, including table information, task descriptions, column exploration strategies, format restrictions, and API-specific query guidelines. During each iteration, a new SQL query is generated using the chat session model and executed via the specified API. If the query returns valid results, the results are processed by rounding numeric columns to two decimal places and handling nested values. The results are checked for uniqueness and format compliance, and if they satisfy these criteria, they are stored in a results table (results_tables). Self-consistency is achieved by verifying if the same result appears in results_tables twice, at which point the process terminates. If errors or invalid data are encountered, the algorithm refines the query by addressing issues such as missing or empty columns, improper handling of nested data, or syntax errors, and retries the process with the updated query. The workflow terminates either when valid results meet self-consistency, when the maximum number of iterations (max_iter) is reached, or when three consecutive errors occur, ensuring incomplete or invalid outputs are discarded. This structured approach improves robustness and accuracy in SQL query execution.